OpenStudy (anonymous):

kayla's mother bought 2 1/2 pounds of berries on monday. kayla's family ate 1/4 of the berries Monday night before she went to bed.

OpenStudy (anonymous):

2 1/2 lb berries 1/4 of the berries are eaten. (1/4) * (2 1/2) = (1/4) * (5/2) = (5/8) of the berries were eaten How many of the berries were leftover? 2 1/2 - 5/8 = 5/4 - 5/8

OpenStudy (anonymous):

If that's what you're asking for, you can solve the last part. You didn't really pose a question.

OpenStudy (anonymous):

u r awesome thnx

OpenStudy (anonymous):

thats exactly what i was asking for but i pushed a button and it posted it

OpenStudy (anonymous):

wait so how many were left over in a fraction?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{5}{4}-\frac{5}{8}\] How do you solve this problem? you try to get the same denominator. \[\frac{5}{4}=\frac{5}{4}*\frac{2}{2}=\frac{10}{8}\] Now you can solve more easily \[\frac{10}{8}-\frac{5}{8}\]

OpenStudy (anonymous):

so it is 5/8 left over?

OpenStudy (anonymous):

2 and 1/2 is actually 5/2 not 5/4. that should be 20/8 - 5/8. My earlier math was wrong

OpenStudy (anonymous):

i still need a little help

OpenStudy (anonymous):

so the leftover of berries is what?